### v1

# Towards A Universal Graph Structural Encoder

Abstract
Recent advancements in large-scale pre-training have shown the potential to learn generalizable representations for downstream tasks. In the graph domain, however, capturing and transferring structural information across different graph domains remains challenging, primarily due to the inherent differences in topological patterns across various contexts. Additionally, most existing models struggle to capture the complexity of rich graph structures, leading to inadequate exploration of the embedding space. To address these challenges, we propose GFSE, a universal graph structural encoder designed to capture transferable structural patterns across diverse domains such as molecular graphs, social networks, and citation networks. GFSE is the first cross-domain graph structural encoder pre-trained with multiple self-supervised learning objectives. Built on a Graph Transformer, GFSE incorporates attention mechanisms informed by graph inductive bias, enabling it to encode intricate multi-level and fine-grained topological features. The pre-trained GFSE produces generic and theoretically expressive positional and structural encoding for graphs, which can be seamlessly integrated with various downstream graph feature encoders, including graph neural networks for vectorized features and Large Language Models for text-attributed graphs. Comprehensive experiments on synthetic and real-world datasets demonstrate GFSE’s capability to significantly enhance the model’s performance while requiring substantially less task-specific fine-tuning. Notably, GFSE achieves state-of-the-art performance in $81.6\%$ evaluated cases, spanning diverse graph models and datasets, highlighting its potential as a powerful and versatile encoder for graph-structured data.
Machine Learning, ICML
[…]
1 Introduction
The success of large-scale pre-trained models in domains such as natural language processing (Achiam et al., 2023; Bubeck et al., 2023; Touvron et al., 2023), computer vision (Radford et al., 2021; Ramesh et al., 2021), audio (Yang et al., 2023; Borsos et al., 2023) has demonstrated the power of learning transferable representations across diverse datasets. However, the development of such models remains challenging in the graph domain. Due to the inherent difference in dataset-specific features, most prior graph pre-training models are specialized for certain areas, such as molecules (Zhang et al., 2020; Sypetkowski et al., 2024), proteins (Nijkamp et al., 2023), and knowledge graphs (Galkin et al., 2023). These specialized models require domain-specific knowledge and suffer from limited transferability to different graph domains.
Graphs exhibit rich structural patterns that are often shared across domains. For example, social networks commonly exhibit small-world properties and community structures, while biological networks reveal recurring motifs and hierarchical modularity. Similarly, citation networks share characteristics like scale-free degree distributions and core-periphery structures. Despite their prevalence, they are often overlooked in favor of domain-specific features, limiting the transferability of existing graph pre-trained models.
As tokens in NLP derive meaning from their context, nodes in graphs gain significance through their structural patterns. Therefore, we propose a paradigm shift that focuses on universal structural attributes—domain-agnostic features that are shared across diverse graph domains. These attributes, such as motifs, node degrees, and hierarchical topologies, serve as the building blocks for transferable graph representations. Developing a graph pre-trained model that can capture these universal structural patterns in a foundational manner, while remaining flexible enough to handle domain-specific adaptation, presents significant challenges. These challenges highlight the need for innovative pre-training strategies that prioritize universal graph encoding over domain-specific features.
Proposed work. To address the challenges of cross-domain pre-training and effectively capturing universal structural encoding, we propose GFSE, a Graph Foundational Structural Encoder, as shown in Figure 1. GFSE is pre-trained across diverse graph domains using multiple self-supervised pre-training tasks, including shortest path distance regression, motif counting, local community detection, and graph-level contrastive learning. Each pre-training task targets a critical and necessary aspect of graph structures, enabling GFSE to capture a comprehensive understanding of graph topology. GFSE employs a Graph Transformer enhanced with biased attention mechanisms. Notably, the relative positional encoding, derived from the random walk matrix, is explicitly integrated into the attention bias term. This design allows GFSE to effectively capture intricate structural dependencies among node pairs during pre-training, ensuring both efficiency and theoretically guaranteed expressiveness. GFSE’s versatility extends to various graph learning scenarios. The pre-trained GFSE can produce generic and expressive Positional and Structural Encodings (PSE) for topological tasks. In feature-enriched contexts, the generated PSE can seamlessly augment vectorized features or integrate with text encoders (e.g., LLMs) for text-attributed graphs, enabling GFSE to serve as a powerful component in graph foundation models.
The contributions of this work are threefold. (1) We propose GFSE, the first cross-domain graph structural encoder pre-trained with four essential self-supervised learning objectives. Extensive experiments show the effectiveness of these pre-training tasks, leading to state-of-the-art performance in $81.6\%$ evaluated cases, spanning diverse graph models and datasets. (2) We provide theoretical justification and empirical results demonstrating GFSE’s ability to generate expressive PSE. (3) GFSE serves as a plug-and-play solution for any graph foundation model to incorporate structural information. By focusing on universal structural patterns, GFSE paves the way for more generalizable and adaptable graph encoding, potentially reducing the need for domain-specific fine-tuning in many applications.
Figure 1: A) GFSE is pre-trained on 8 datasets from 6 different domains. Pre-training tasks include reconstruction (shortest path distance regression, motif counting) and contrastive learning (community detection, graph contrastive learning). B) GFSE generates generic and expressive Positional and Structural Encoding (PSE) to tackle topological graph tasks. GFSE can also be seamlessly integrated into downstream feature encoders for feature-enriched tasks by concatenating with initial vectorized features or prepending the generated PSE to the textual prompt as a soft token.
2 Related Work
Graph Pre-training. Graph self-supervised learning approaches are typically pre-training graph models, e.g., GNNs or Graph Transformers, on a massive amount of labeled graphs with inherent features by reconstructing the structures or masked attributes (Cui et al., 2020; Hou et al., 2022; Kipf & Welling, 2016b; Hu et al., 2020b; Wang et al., 2017; Xia et al., 2024; Xia & Huang, 2024; Zhao et al., 2024b; Mizera et al., 2024). Some works also utilize contrastive learning to enhance node and graph-level representation learning (Han et al., 2022; Hassani & Khasahmadi, 2020; Velickovic et al., 2019; Hu et al., 2019; Lee et al., 2022; Li et al., 2021; Lu et al., 2021; Sun et al., 2019, 2021; Xu et al., 2021a; Galkin et al., 2023; Zhao et al., 2024a). These methods, while effective in certain domains, exhibit limited generalizability across different graph domains due to their tailored design for specific types of data. Additionally, there have been graph prompt techniques (Huang et al., 2024; Fang et al., 2024) that can be used to enhance model adaptation over graphs. There have also been some attempts at cross-domain graph pre-training models (Qiu et al., 2020; Davies et al., 2023). Unfortunately, all these models rely on one singular pre-training task (i.e., contrastive learning), and usually fail to capture fine-grained structural features at node level or edge level (Mao et al., 2024).
Graph Language Models. With the success of foundation models in NLP, recent efforts also harness LLMs to develop domain-specific graph foundation models by flattening graph structures and associated texts into prompts (Chen et al., 2024a; Tang et al., 2023; Ye et al., 2023; Qian et al., 2023; Zhao et al., 2023; Guo et al., 2023; Chen et al., 2024b; Liu et al., 2023a; Kong et al., 2024; Chen et al., 2024c; Fan et al., 2024; Zhang et al., 2024; Li et al., 2024). Nevertheless, recent studies show that LLM demonstrates an unsatisfying ability to reason and understand complicated structures within graph (Fatemi et al., 2023; Zhao et al., 2023; Wang et al., 2024). In this work, we focus on developing an expressive graph model dedicated to encoding the rich topological information, without relying on associated text. Our approach complements LLMs on text-attributed graphs, serving as a universal structural encoder.
Positional and Structural Encoding (PSE). Traditional PSEs include hand-crafted features such as Laplacian PE (Davies et al., 2024; Kreuzer et al., 2021; Beaini et al., 2021; Wang et al., 2022), shortest-path distance (Li et al., 2020; Ying et al., 2021), kernel distance (Mialon et al., 2021), random-walk encoding (Ma et al., 2023; Dwivedi et al., 2021; Brüel-Gabrielsson et al., 2022; Rampášek et al., 2022), node degree centrality (Ying et al., 2021), etc. Some studies have introduced specialized networks designed to adaptively learn PSE to enhance performance (Kreuzer et al., 2021; Dwivedi et al., 2021; Chen et al., 2022; Lim et al., 2022). GPSE (Liu et al., 2023b) proposes to pre-train a positional and structural encoder on domain-specific graphs to generate PSE. However, GPSE still suffers from limited transferability and expressiveness across other domains, due to its simplistic pre-training backbone and randomized node features. Finding the most effective and versatile PSE remains an open challenge that requires further innovation.
3 Proposed Method
As shown in Figure 1, we collect graph pre-training datasets from six different domains, including molecules, proteins, social networks, images, product networks, and academic networks. GFSE utilizes a transformer-based architecture with biased attention to incorporate relative inductive bias within graph structures (Sec. 3.1). GFSE is pre-trained with four challenging self-supervision tasks simultaneously, each designed to enhance a crucial aspect of structural awareness and promote encoding quality (Sec. 3.2). GFSE generates expressive positional and structural encoding (PSE) for topological tasks. The generated PSE can also be seamlessly integrated into graphs with vectorized features or textual features, enhancing the downstream performance (Sec. 3.3).
3.1 Architecture
Previous work (Liu et al., 2023b) uses randomized features to replace initial node features. However, it leads to poor generalizability across different domains. In this work, we propose to use both absolute and relative random-walk positional encoding as the initial features. Formally, let $G(V,E)$ represent an input graph, where $V$ and $E$ denote the set of nodes and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ indicates the adjacency matrix, where $N$ is the number of nodes, and $\mathbf{D}$ is the degree matrix. Random Walk matrix is defined as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{M}_{i,j}$ indicates the transition probability from the $i$-th node to the $j$-th node. Following previous works (Ma et al., 2023), we calculate the $d$-dimensional encoding for each node and node pairs as follows.
$\mathbf{P}_{i}$ $=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,i}$ (1)
$\mathbf{R}_{i,j}$ $=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,j}$ (2)
$\mathbf{P}\in\mathbb{R}^{N\times d}$ and $\mathbf{R}\in\mathbb{R}^{N\times N\times d}$ are used as the initial node features and edge features in GFSE.
Pre-training Backbone. GFSE is built on a GPS architecture (Rampášek et al., 2022) for pre-training, due to its scalability and generalizability. Each GPS layer contains local message passing and global attention modules to capture neighbor and long-range information. In the $\ell$-th layer, the node encoding $\mathbf{P}^{\ell}$ and relative edge encoding $\mathbf{R}^{\ell}$ are fed into message passing layers ($\operatorname{MPNN}$) and Biased Attention Module ($\operatorname{BiasAttn}$) parallelly. The node encoding $\mathbf{P}^{\ell+1}$ is updated by aggregating local and global information.
$\vspace{-0.5cm}\mathbf{P}_{M}^{\ell+1},\mathbf{R}^{\ell+1}$ $=\operatorname{MPNN}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell},\mathbf{A}\right)$ (3)
$\mathbf{P}_{T}^{\ell+1}$ $=\operatorname{BiasAttn}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell}\right)$ (4)
$\mathbf{P}^{\ell+1}$ $=\operatorname{MLP}^{\ell}\left(\mathbf{P}_{M}^{\ell+1}+\mathbf{P}_{T}^{\ell+1}\right)$ (5)
Attention Bias. The global attention in the original GPS framework does not account for relative edge encoding while leaving them entirely for the message-passing layers. However, incorporating relative edge encoding in global attention is crucial for capturing long-range dependencies, as the receptive field of message-passing layers is inherently constrained by their depth. To build a theoretically more powerful GPS, we explicitly incorporate relative edge encoding into global attention, where the attention weight between the $i$-th and the $j$-th nodes is computed by $a_{i,j}^{\prime}=\operatorname{SoftMax}(a_{i,j}+\operatorname{Linear}(\mathbf{R}_{i,j}))$, where $\operatorname{Linear}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ indicates a linear layer that maps the $d$-dimensional relative encoding to a scalar. $a_{i,j}$ denotes the original attention weight computed by scaled-dot self-attention on the node encoding $\mathbf{P}^{\ell}$ in each GPS layer. We refer to Appendix C.3 (Figure 5) for biased attention details.
3.2 Self-supervised Pre-training Tasks
GFSE is pre-trained with reconstruction and contrastive structural tasks at multiple context levels. Each task highlights a specific structural aspect, thereby augmenting the model’s expressiveness and capability to capture complex graph structures. Let $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$ represent the output after $L$ GPS layers. We decode $\mathbf{P}^{L}$ with independent MLP heads for each pre-training task.
[…]
Motif Counting enables the model to better identify each node’s role in the surrounding subgraphs. To improve expressiveness, we include a variety of small motifs, called graphlets (Pržulj et al., 2004; Pržulj, 2007; Bouritsas et al., 2022), with different numbers of vertices, which are beyond usual types like stars, paths, cycles, and cliques. We refer to Appendix B.2 for more technical details. Let $Y_{i}^{m}\in\mathbb{Z}^{k}$ denote the motif label that indicates the number of certain motifs surrounding node $i$, where $k$ is the number of graphlet types. The loss is formulated as $\mathcal{L}_{\text{MC}}=\frac{1}{|V|}\sum_{i\in V}\|h_{\text{MC}}(\mathbf{P}^{L}_{i})-Y_{i}^{m}\|^{2}$, where $h_{\text{MC}}:\mathbb{R}^{d}_{e}\rightarrow\mathbb{R}^{k}$ is the task-specific head for motif counting.
Community Detection aims to identify densely connected subgraphs, where nodes within a community are more closely linked to each other than to nodes outside the community. Such community structures are ubiquitous in various real-world networks, e.g., social networks, and transportation systems. We employ the Louvain Community Detection Algorithm (Blondel et al., 2008) to extract the community structure from pre-training graphs, which clusters nodes into communities based solely on graph topology without node features. We approach this task in a contrastive learning manner by minimizing the embedding distances between intra-community nodes while maximizing the distance between inter-community nodes.
$\mathcal{L}_{\text{CD}}=$ $\sum_{i\in V}\sum_{j\in V}Y_{i,j}^{c}(1-\operatorname{sim}(i,j))$
$+(1-Y_{i,j}^{c})\max(0,\epsilon-(1-\operatorname{sim}(i,j))),$
where the similarity score $\operatorname{sim}(i,j)$ is calculated by $\operatorname{sim}(i,j)=\frac{\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{j}}{||\boldsymbol{z}_{i}||\cdot||\boldsymbol{z}_{j}||}$ and $\boldsymbol{z}_{i}=h_{\text{CD}}(\mathbf{P}^{L}_{i})$ with a head $h_{\text{CD}}$. $\epsilon$ is a margin hyperparameter. $Y_{i,j}^{c}$ is a binary label that indicates if the $i$-th node and the $j$-th node are in the same community. Through $\mathcal{L}_{\text{CD}}$, GFSE learns to discern community boundaries and distinguish nodes based on local community structures.
Graph Contrastive Learning aims to distinguish graphs from different datasets. The motivation is rooted in the observation that similar structural characteristics in different domains may exhibit distinct meanings. For example, a subgraph representing a protein interaction network might correspond to a specific functional module, whereas a similar structure in a social network could represent a tightly-knit community or interest group. Therefore, GFSE distinguishes graphs from different datasets, in addition to performing structural pre-training tasks within a single graph. The loss for this task is computed by
$\mathcal{L}_{\text{GCL}}=-\log\frac{\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}}\right)/\tau\right)}{\sum_{k=1}^{K}\mathbbm{1}_{[G_{k}\nsim G_{i}]}\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{k}}\right)/\tau\right)},$
where $\tau$ is the temperature, $G_{i}$ and $G_{j}$ are from the same dataset, $\boldsymbol{z}_{G_{i}}=\mathrm{GlobalPool}(h_{\text{GCL}}(\mathbf{P}^{L}_{G_{i}}))$ is the output of the global pooling applied to the final layer’s representation $\mathbf{P}^{L}_{G_{i}}$ for the graph $G_{i}$, $K$ is the number of negative samples, and $\mathbbm{1}_{[G_{k}\nsim G_{i}]}$ is an indicator function that determines whether graphs $G_{k}$ and $G_{i}$ originate from different datasets.
Multi-task Selection and Loss Balance. Each pre-training task targets a different structural aspect, enabling GFSE to capture a comprehensive understanding of graph topology. For instance, the shortest path distance regression task emphasizes global connectivity within graphs, while motif counting focuses on the occurrence of local subgraph patterns. Although other pre-training tasks could be considered, the chosen tasks provide a balanced and effective combination, capable of capturing both local and global structural information at varying granularities without introducing unnecessary complexity. Since the loss scale and difficulty vary significantly across tasks, we introduce task-specific uncertainty (Kendall et al., 2018), which is learnable during pre-training to unify the scales of all losses. Task-specific uncertainty is used to automatically balance different pre-training losses, i.e., $\mathcal{L}_{\text{SPD}},\mathcal{L}_{\text{MC}},\mathcal{L}_{\text{CD}}$, and $\mathcal{L}_{\text{GCL}}$ (see Appendix C.4 for more details). Moreover, the evolution of uncertainty values provides insights into each task’s contribution to the overall pre-training process.
3.3 Combination with Downstream Feature Encoder
Application on Graphs with Vectorized Features. GFSE can be readily employed to generate expressive PSE for various graph applications. Let $\mathbf{X}^{0}\in\mathbb{R}^{N\times d_{x}}$ denote the initial node features for a given graph with $N$ nodes and $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$ denote PSE generated by GFSE, where $d_{x}$ and $d_{e}$ are dimensions of node features and PSE, respectively. $\mathbf{P}^{L}$ can then be concatenated with the initial node features $\mathbf{X}^{0}$ to create a new feature matrix $\mathbf{X}^{new}=[\mathbf{X}^{0}\|\mathbf{P}^{L}]\in\mathbb{R}^{N\times(d_{x}+d_{e})}$, which augments the node features with structural information. This structure-enriched feature $\mathbf{X}^{new}$ can subsequently be fed into downstream graph models, such as graph neural networks or graph transformers, enhancing their performance on various tasks. For large-scale graphs, where computing PSE for the entire graph may be computationally prohibitive, we thereby sample the neighborhood structure around each node and compute the PSE for these localized subgraphs, which is efficiently parallelized, enabling scalable and efficient generation of PSE for large graphs.
Application on Text-attributed Graphs. Language models are typically employed to process the text-attributed graphs, where GFSE can be seamlessly applied to incorporate structural information within graphs. Given the generated $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$, an MLP is employed to project $\mathbf{P}^{L}$ into the embedding space of the language model. This projected PSE is then prepended as a soft token to the associated text, effectively incorporating the graph’s structural information into the model input. Subsequently, these structure-enriched tokens are fed into downstream large language models (LLMs), enhancing their performance on graph-related tasks. The process, involving training a lightweight MLP and fine-tuning LLM with Parameter-Efficient Fine-Tuning (PEFT) techniques such as LoRA (Hu et al., 2021), makes it scalable and efficient for large-scale text-attributed graph applications.
3.4 Expressive Power of GFSE
We show that GFSE can generate highly expressive PSE by incorporating relative edge encoding into the attention computation in the Graph Transformer backbone. Specifically, we employ the Structural Encoding enhanced Global Weisfeiler-Lehman test (SEG-WL) (Zhu et al., 2023), a generalized WL test that incorporates relative structural encoding into the isomorphism algorithm, to characterize the expressiveness of GFSE. For an input graph $G(V,E)$ with node set $V$ and edge set $E$, let $f_{P}:V\rightarrow\mathcal{X}$ and $f_{R}:V\times V\rightarrow\mathcal{X}$ indicate the node-level and edge-level structural encoding, respectively. Different from traditional WL test, SEG-WL updates the node labels at the $t$-th iteration by $g_{t}(v)=\text{hash}\left(\left\{\left\{\left(g_{t-1}(u),f_{R}(v,u)\right):u\in V\right\}\right\}\right)$ and $g_{0}(v)=\text{hash}(f_{P}(v))$. SEG-WL can be viewed as a high-level abstraction of the learning paradigm of our pre-training architecture with biased attention (Eq. 3), where relative structural encoding between any two nodes is considered for updating node representations (Zhu et al., 2023). Let RW($d$)-SEG-WL denote the case that $f_{P}$ and $f_{R}$ are determined by $\mathbf{P}$ and $\mathbf{R}$ with $d$ dimension, i.e., $f_{P}(v_{i})=\mathbf{P}_{i}\in\mathbb{R}^{d}$ and $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ for the $i$-th and $j$-th nodes. We have the following propositions.
Proposition 3.1.
RW($d$)-SEG-WL ($d\geq 3$) is strictly more expressive than 1-WL in testing non-isomorphic graphs.
Proposition 3.2.
There exist pairs of graphs that RW($d$)-SEG-WL can distinguish, but 3-WL can not.
The theoretical proof and empirical verification are given in Appendix D. RW-SEG-WL is able to distinguish all low-order graphs with orders equal to or less than 8 and successfully distinguishes most strongly regular graphs where 3-WL fails to distinguish. RW-SEG-WL stands as an expressivity upper bound of our proposed GFSE. The pre-training tasks are meticulously designed to push GFSE towards achieving the upper bound established by RW-SEG-WL. These pre-training tasks optimize both node-level and edge-level structural encoding, progressively refining the effectiveness of the model in generating expressive encoding.
3.5 Computational Complexity
GFSE’s complexity comprises two parts: pre-computation of self-supervision labels and pre-training. Computing the shortest path distance between node pairs results in the time complexity of $\mathcal{O}(|E|+|V|\log|V|)$ with node set $V$ and edge set $E$. A brute-force implementation of the subgraph isomorphism counting of fixed size $t$ is $\mathcal{O}(|V|^{t})$. We consider the graphlets with at most 5 nodes. One can also choose special graphlet types, e.g., paths, cycles, and triangles, which can be efficiently enumerated (Giscard et al., 2019). Approximating and scalable algorithms can be further used to accelerate this pre-processing step (Fu et al., 2024; Ying et al., 2020; Pashanasangi & Seshadhri, 2020). For pre-training, the complexity is $\mathcal{O}(|V|^{2})$ for full attention computation and $\mathcal{O}(d|V||E|)$ for initial encoding computation. See runtime evaluations in Appendix F.2.
4 Experiments
We first evaluate the pre-training performance in Sec. 4.2 and empirically assess the expressiveness of GFSE on synthetic datasets (Sec. 4.3). We then evaluate GFSE in a wide range of downstream graph learning tasks in Sec. 4.4. Specifically, we conduct experiments with LLMs on text-attributed graphs in Sec. 4.5.
4.1 Pre-Training Setup
Dataset. We utilize diverse datasets for pretraining, ensuring a broad spectrum of graph structures and scales, including MolPCBA, MolHIV, MNIST, peptides, ogbn-proteins, Pokec, ogbn-arxiv and ogbn-product (Wu et al., 2018; Bhatia et al., 2016; Mikolov et al., 2013; Szklarczyk et al., 2019; Chiang et al., 2019; Takac & Zabovsky, 2012; Dwivedi et al., 2023, 2022). These datasets cover several real-world graph domains, such as social networks, academic networks, etc. We refer to Table 6 in Appendix C.1 for detailed dataset statistics. For large-scale graphs, we first partition them into sets of subgraphs by the METIS algorithm (Karypis & Kumar, 1997) to handle scalability issues. Training samples from different datasets are mixed and randomly shuffled to form a large-scale pre-training dataset.
Training Setting. The pre-training stage is conducted on the standard train/validation/test splits of the pre-training datasets. The initial encoding dimension $d$ is set as 8. We adopt GIN (Xu et al., 2018) as the message-passing layer in the GPS and adopt 8 GPS layers with 8 heads and 128 hidden dimensions. The output dimension is 64 by default. We use Adam as the optimizer with an initial learning rate of $0.001$ and the batch size is 256. The maximum training epochs is 100. An early stopping strategy is used to mitigate overfitting. The pre-training is implemented on the NVIDIA A40 48GB GPU. We refer to Appendix C.5 for more details. The code is available at https://github.com/Graph-and-Geometric-Learning/GFSE.
4.2 Pre-training Evaluation
GFSE is pre-trained with four self-supervised learning tasks. We iteratively change the message passing layers (e.g., GatedGCN (Bresson & Laurent, 2017), GCN (Kipf & Welling, 2016a) and GIN (Xu et al., 2018)) and replace the biased attention with traditional self-attention in the default GFSE architecture. We evaluate the pre-training performance with different architectures as shown in Figure 2. Accuracy is used to measure community detection and graph contrastive learning tasks, indicating the proportion of node (graph) pairs that are correctly predicted. MSE and MAE are used for shortest path distance and motif counting tasks. See more details in Appendix C.2.
Figure 2: Pre-training performance with different architectures. TF is the abbreviation of transformer.
We observe a consistent performance boost when combining biased attention with GIN, which serves as the default configuration of GFSE. The pre-trained GFSE is able to generate effective PSE, which can reconstruct other pre-defined PSE, such as LapPE (Lim et al., 2022) and ElstaticPE (Kreuzer et al., 2021). We provide experimental results in Appendix D.3. The ability to reconstruct various types of PSE, without taking them as training objectives explicitly, demonstrates the sufficiency and effectiveness of the chosen tasks in pre-training the model for expressive and generalizable graph representations.
4.3 Expressiveness Evaluation
We empirically evaluate the structure-awareness of the positional and structural encoding (PSE) generated by GFSE on Triangle (Knyazev et al., 2019), Pattern, and Cluster (Dwivedi et al., 2023) , three datasets that require discerning intricate graph topologies. We evaluate the performance boost brought by GFSE in comparison to traditional RWSE and LapPE. We test on various downstream graph learning models, including MLP, GIN (Xu et al., 2018), transformer (Vaswani et al., 2017) and GPS (Rampášek et al., 2022). We further compare with the learning-based PSE approach: GPSE (Liu et al., 2023b). See Appendix A for more baseline details. As shown in Table 1, GFSE generates expressive and robust PSE that consistently improves the base model’s performance, whereas other structural encodings exhibit considerable variation across different datasets or different base models. Notably, the performance boost brought by GFSE is particularly significant for the Transformer on the Triangle dataset. This demonstrates that GFSE has a stronger enhancement effect on models that originally lack structural bias.
[…]
Triangle-S Triangle-L Pattern Cluster
MLP+RWSE 98.22 11.88 50.53 20.96
MLP+LapPE 98.60 12.62 50.53 20.96
MLP+GPSE 52.80 17.42 55.66 20.96
MLP+GFSE 98.71 25.54 57.79 21.28
GIN 99.68 42.58 85.58 60.84
GIN+RWSE 99.70 40.78 85.34 61.30
GIN+LapPE 99.74 42.48 85.45 61.83
GIN+GPSE 99.32 25.32 85.19 61.95
GIN+GFSE 99.72 43.84 85.58 63.49
Transformer (TF) for Triangle and GPS for Others
TF / GPS 21.68 23.58 86.63 77.76
TF / GPS+RWSE 35.96 11.38 86.68 77.72
TF / GPS+LapPE 35.96 12.44 86.54 77.76
TF / GPS+GPSE 62.04 24.64 85.58 77.80
TF / GPS+GFSE 92.82 30.15 87.98 77.86
4.4 Downstream Evaluation
4.4.1 Evaluation on Pre-training Dataset
Dataset and Baseline. We conduct a comprehensive evaluation of GFSE on real-world graph datasets. We adhere to the experimental setting and hyper-parameters established by previous works (Rampášek et al., 2022) to implement base models. We augment the initial node features with GFSE and evaluate the performance in downstream tasks, comparing it against established structural encoding methods including RWSE, LapPE, and GPSE (Liu et al., 2023b).
[…]
Table 2: Performance on MolPCBA, ZINC (subset), Peptides-func and Peptides-struct.
MolPCBA ZINC Peptides-func Peptides-struct
AP $\uparrow$ MAE $\downarrow$ AP $\uparrow$ MAE $\downarrow$
GCN $0.2424_{\pm 0.0034}$ $0.3670_{\pm 0.0110}$ $0.5930_{\pm 0.0023}$ $0.3496_{\pm 0.0013}$
GCN+LapPE $0.2417_{\pm 0.0047}$ $0.2052_{\pm 0.0132}$ $0.6021_{\pm 0.0051}$ $0.2688_{\pm 0.0027}$
GCN+RWSE $0.2438_{\pm 0.0028}$ $0.1741_{\pm 0.0528}$ $0.5827_{\pm 0.0046}$ $0.3270_{\pm 0.0019}$
GCN+GPSE $0.1958_{\pm 0.0074}$ $\mathbf{0.1218}_{\pm 0.0613}$ $0.5959_{\pm 0.0034}$ $0.2710_{\pm 0.0041}$
GCN+GFSE $\mathbf{0.2477}_{\pm 0.0021}$ $0.1237_{\pm 0.0428}$ $\mathbf{0.6131}_{\pm 0.0074}$ $\mathbf{0.2513}_{\pm 0.0054}$
GIN $0.2703_{\pm 0.0023}$ $0.5260_{\pm 0.0510}$ $0.5498_{\pm 0.0079}$ $0.3547_{\pm 0.0045}$
GIN+LapPE $0.2701_{\pm 0.0013}$ $0.2203_{\pm 0.0386}$ $0.5323_{\pm 0.0083}$ $0.2650_{\pm 0.0041}$
GIN+RWSE $0.2781_{\pm 0.0031}$ $0.1731_{\pm 0.0614}$ $0.5410_{\pm 0.0068}$ $0.3282_{\pm 0.0037}$
GIN+GPSE $0.2765_{\pm 0.0073}$ $0.2162_{\pm 0.0429}$ $0.5389_{\pm 0.0094}$ $\mathbf{0.2581}_{\pm 0.0046}$
GIN+GFSE $\mathbf{0.2839}_{\pm 0.0046}$ $\mathbf{0.1689}_{\pm 0.0524}$ $\mathbf{0.5532}_{\pm 0.0103}$ $0.2674_{\pm 0.0039}$
Transformer (TF) $0.0808_{\pm 0.0117}$ $0.6943_{\pm 0.0328}$ $0.4800_{\pm 0.0076}$ $0.4192_{\pm 0.0028}$
TF+LapPE $0.1784_{\pm 0.0329}$ $0.5101_{\pm 0.0724}$ $0.6307_{\pm 0.0091}$ $0.2514_{\pm 0.0031}$
TF+RWSE $0.2083_{\pm 0.0674}$ $0.2193_{\pm 0.0640}$ $0.6326_{\pm 0.0028}$ $0.3344_{\pm 0.0028}$
TF+GPSE $0.2040_{\pm 0.0531}$ $0.1883_{\pm 0.0263}$ $0.6534_{\pm 0.0041}$ $0.2479_{\pm 0.0068}$
TF+GFSE $\mathbf{0.2376}_{\pm 0.0342}$ $\mathbf{0.1548}_{\pm 0.0796}$ $\mathbf{0.6642}_{\pm 0.0025}$ $\mathbf{0.2436}_{\pm 0.0071}$
GPS $0.2869_{\pm 0.0045}$ $0.1182_{\pm 0.0049}$ $0.6535_{\pm 0.0041}$ $0.2500_{\pm 0.0012}$
GPS+LapPE $\mathbf{0.2939}_{\pm 0.0016}$ $0.1078_{\pm 0.0084}$ $0.6494_{\pm 0.0037}$ $0.2501_{\pm 0.0026}$
GPS+RWSE $0.2907_{\pm 0.0028}$ $0.0700_{\pm 0.0040}$ $0.6603_{\pm 0.0085}$ $0.2739_{\pm 0.0063}$
GPS+GPSE $0.2911_{\pm 0.0036}$ $0.0648_{\pm 0.0030}$ $0.6688_{\pm 0.0151}$ $\mathbf{0.2464}_{\pm 0.0025}$
GPS+GFSE $0.2916_{\pm 0.0061}$ $\mathbf{0.0613}_{\pm 0.0026}$ $\mathbf{0.6874}_{\pm 0.0120}$ $0.2474_{\pm 0.0051}$
[…]
4.4.2 Evaluation on Molecular Graphs
Settings and Methods. We use the molecular datasets namely Tox21, Sider, BBBP, and MUV (Hu et al., 2020a) as a downstream task. We evaluate the effectiveness of GFSE under two settings: (1) In the training from scratch setting, we directly concatenate GFSE’s PSE with the raw node features to create new input features. This augmented representation is then fed into a randomly initialized model from the beginning of training. We take GINE (Xu et al., 2018) and GPS (Rampášek et al., 2022) as our backbone. (2) In the fine-tuning setting, we assess GFSE’s ability to enhance pre-trained models by concatenating the node encodings obtained from a pre-trained model with the GFSE’s PSE. The concatenated features are then fed into the final read-out layers for prediction. During fine-tuning, the parameters of the pre-trained model and read-out layers are updated. We select GraphMAE (Hou et al., 2022) and MoleBERT (Xia et al., 2022) as the pre-trained backbones and compare with other baselines without structural encoding, namely SSP (Hu et al., 2019), GraphLoG (Xu et al., 2021b), GraphCL (You et al., 2020). Refer to Appendix E.1 for more implementation details.
Results. Experimental results are shown in Table. 3. For training the models from scratch, on both GINE and GPS, PSE consistently improves model performance, achieving better results than all the other structural feature augmentation methods. As to fine-tuning, GFSE significantly boosts the performance and achieves state-of-the-art performance.
Table 3: Test ROC-AUC(%) performance on molecular datasets. The best results for each dataset are bolded.
Tox21 Sider BBBP MUV
SSP 76.8 $\pm$ 0.8 61.7 $\pm$ 0.8 67.9 $\pm$ 0.9 79.8 $\pm$ 1.6
GraphCL 75.7 $\pm$ 0.5 60.8 $\pm$ 0.7 69.5 $\pm$ 0.5 74.5 $\pm$ 1.3
GraphLoG 75.4 $\pm$ 0.9 61.2 $\pm$ 1.1 72.5 $\pm$ 0.8 76.0 $\pm$ 1.1
Train-from-scratch Models
GINE 74.5 $\pm$ 0.4 58.6 $\pm$ 0.1 67.7 $\pm$ 0.7 74.8 $\pm$ 0.6
GINE+RWSE 75.3 $\pm$ 0.2 58.4 $\pm$ 1.8 66.7 $\pm$ 1.4 76.4 $\pm$ 0.8
GINE+LapPE 77.6 $\pm$ 0.8 57.2 $\pm$ 1.1 65.8 $\pm$ 0.3 77.0 $\pm$ 0.8
GINE+GPSE 74.9 $\pm$ 0.4 60.1 $\pm$ 0.8 66.4 $\pm$ 0.1 75.8 $\pm$ 1.3
GINE+GFSE 75.5 $\pm$ 0.7 60.9 $\pm$ 0.5 69.1 $\pm$ 1.3 77.7 $\pm$ 1.2
[…]
GPS+RWSE 74.6 $\pm$ 1.3 56.4 $\pm$ 0.6 67.9 $\pm$ 1.0 69.7 $\pm$ 0.6
GPS+LapPE 74.8 $\pm$ 1.1 60.5 $\pm$ 0.6 67.9 $\pm$ 0.6 70.1 $\pm$ 2.2
GPS+GPSE 75.1 $\pm$ 0.7 56.6 $\pm$ 1.7 67.8 $\pm$ 0.7 68.3 $\pm$ 0.1
GPS+GFSE 76.3 $\pm$ 1.4 61.8 $\pm$ 0.5 68.0 $\pm$ 0.5 73.6 $\pm$ 0.5
Pre-trained Models
GraphMAE 75.4 $\pm$ 0.4 59.8 $\pm$ 0.5 69.5 $\pm$ 1.6 76.3 $\pm$ 2.4
GraphMAE+RWSE 76.3 $\pm$ 0.5 60.5 $\pm$ 0.8 66.4 $\pm$ 3.7 77.7 $\pm$ 1.5
GraphMAE+GFSE 75.9 $\pm$ 0.9 62.1 $\pm$ 0.8 70.5 $\pm$ 1.4 78.1 $\pm$ 1.3
[…]
MoleBERT+RWSE 77.8 $\pm$ 0.7 63.1 $\pm$ 0.6 66.5 $\pm$ 2.1 80.4 $\pm$ 1.3
MoleBERT+GFSE 78.0 $\pm$ 0.4 63.1 $\pm$ 0.7 68.9 $\pm$ 2.1 80.5 $\pm$ 2.0
4.5 Integration with Large Language Models
Settings and Methods. We perform experiments on citation networks and e-commerce networks, including Arxiv (Hu et al., 2020a), Cora (Yang et al., 2016) and Amazon-product (He & McAuley, 2016; McAuley et al., 2015), which are text-attributed graphs with detailed descriptions for each node (e.g., product item). Edges indicate co-authored or co-purchased relations between two nodes. We employ a lightweight MLP to align the PSE generated by GFSE with the language model’s embedding space, which ensures seamless integration of structural information into the language model. We concatenate the textual description of a central node with those of its one-hop neighbors and prepend the PSE as a soft token, followed by a special graph token at the end. This combined sequence is then encoded by a language model. The hidden embedding of the special graph token is used as the representation for the central node. Following the previous setting (Zhu et al., 2024), we compute the cosine similarity between the representations of node pairs as the edge likelihood. We train the MLP and fine-tune the language model with LoRA (Hu et al., 2021) using a contrastive loss (Hadsell et al., 2006). More evaluation details can be found in Appendix E.2.
Table 4: Evaluation on text-attributed graphs
Cloth Home Sport Arxiv Cora
Hit@1 MRR Hit@1 MRR Hit@1 MRR ACC ACC
InstructGLM 76.23 82.60 79.82 85.93 62.50 73.25 74.94 79.95
Finetuned LLaMA 74.73 82.87 78.93 86.07 62.52 75.77 75.70 87.08
   + RRWP 76.08 84.11 74.64 85.24 62.71 75.20 76.03 90.27
   + GraphSAGE 76.22 84.16 73.74 81.66 62.26 75.36 76.13 90.83
   + GFSE (Ours) 76.84 84.68 79.85 86.77 64.79 76.24 76.10 91.26
Results. Evaluation results on text-attributed graphs are reported in Table 4. We select InstructGLM (Ye et al., 2023) as a baseline, which has been fine-tuned on graph domains without structural information infusion. Additionally, we include comparisons against RRWP (Ma et al., 2023), a random-walk based PSE, and GraphSAGE (Hamilton et al., 2017), which was trained from scratch as a PSE encoder with LLM finetuning together. Finetuned LLaMA refers to the LLaMA model (Touvron et al., 2023) fine-tuned without any PSE. As seen from the table, GFSE, which is pre-trained on cross-domain graph data as a structural encoder, outperforms other methods in general. In particular, GFSE achieves an average performance boost of $2.66\%$ over InstructGLM, $2.02\%$ over GraphSAGE encoder and $1.53$ over RRWP across the five datasets. These gains demonstrate the effectiveness of GFSE in language-based graph tasks. See more discussions in Appendix F.1.
4.6 More Experimental Analysis
Table 5: Ablation studies on the pre-training tasks and model architecture of GFSE. Best results are shown in bold.
ZINC CIFAR10 Arxiv
MAE $\downarrow$ ACC $\uparrow$ ACC $\uparrow$
GPS w/o PE 0.1182 71.49 70.68
Augment by GFSE 0.0613 74.11 72.30
[…]
   w/o Community Detection 0.0637 72.38 70.34
   w/o Motif Counting 0.0731 73.02 71.27
   w/o Shortest Path Distance Regression 0.1074 71.58 72.06
   w/o Graph Contrastive Learning 0.0856 73.02 72.13
[…]
   GIN+Traditional Attention 0.0872 73.13 71.85
   Biased Attention Only 0.1137 70.97 71.33
   GIN Only 0.0640 72.31 72.34
Ablation on Tasks and Model Architecture. We analyze the sensitivity of each pre-training task and model architecture w.r.t. the performance boost in downstream tasks. The results are illustrated in Table 4.6. Firstly, we iteratively remove one pre-training task and follow the same setting to pre-train and evaluate GFSE. We observe that each removal results in a discernible reduction in downstream performance. Notably, the shortest path distance task is particularly critical for the ZINC and CIFAR10 datasets, while community detection is significant for academic datasets.
We further conduct ablation studies on the main components of GFSE. Specifically, we use traditional attention to replace biased attention as a baseline and remove GIN or attention modules respectively. We notice that all the above pre-training architecture variants lead to performance degradation. The hybrid approach outperforms both attention-only and GIN-only setups, suggesting that integrating sophisticated attention mechanisms can compensate for the absence of global information in local message-passing layers.
Task-specific Uncertainty. Figure 3(a) illustrates the trajectory of task uncertainty ($\sigma^{2}$) across different pre-training tasks w.r.t. pre-training epochs. Higher values of $\sigma^{2}$ reduce the respective task’s contribution to the overall training loss. We observe that all tasks show a sharp decline in uncertainty during the pre-training process. Notably, motif counting maintains a lower uncertainty throughout the training process compared to other tasks, suggesting that it might be inherently more straightforward for the model to optimize or more integral to the model’s overall learning structure.
Figure 3: (a) Learning task uncertainty ($\sigma^{2}$) w.r.t. pre-training. (b) Performance of GFSE pre-trained on different data sizes. Evaluation datasets are MolPCBA and Arxiv. The base model is GPS.
Impact of Pre-training Data Size. Figure 3(b) illustrates the GFSE’s performance when pre-trained on varying ratios of data, revealing several key insights. (1) The model performance declines as the pre-training data size decreases, highlighting the importance of a diverse and extensive pre-training dataset for enhancing the model effectiveness. (2) Pre-training is a critical factor for achieving superior and expressive graph representations, as evidenced by the consistent outperformance of pre-trained GFSE over its train-from-scratch counterpart. (3) Cross-domain pre-training proves to be important for domain generalization. GFSE pre-trained on a single dataset struggles to generalize to different domains (e.g., from MolPCBA to Arxiv). This indicates that models trained on a narrow domain may overfit to the specific characteristics, limiting their generalization to diverse downstream tasks. In contrast, cross-domain pre-training enables the model to learn more transferable and robust representations, which are essential for achieving strong performance across a wide range of applications.
5 Conclusion
GFSE is a significant achievement in universal graph encoder, leveraging multiple self-supervised learning objectives to capture comprehensive structural information from diverse graph domains. By integrating relative positional encoding within a Graph Transformer, GFSE provides a robust framework for generating expressive positional and structural encodings. Extensive experiments on synthetic and real-world datasets validate GFSE’s ability to enhance the performance of various graph feature encoders, broadening its applicability across numerous graph-related tasks. Building upon the promising results of GFSE, a promising future direction could involve dynamic graph structures in the model pre-training, allowing it to adapt to evolving graph topologies over time.
Appendix A Positional and Structural Encoding
Random Walk Structural Encoding. Random walk structural encoding (RWSE) (Dwivedi et al., 2021; Rampášek et al., 2022) uses landing probabilities derived from random walks of varying lengths, starting from each node in the graph to capture both structural and positional relationships among nodes. Formally, let $G(V,E)$ represent an input graph, where $V$ and $E$ denote the set of $N$ nodes and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ indicates the adjacency matrix and $\mathbf{D}$ is the degree matrix. Random Walk matrix is defined as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{M}_{i,j}$ indicates the transition probability from the $i$-th node to the $j$-th node. The RWSE with $d$ steps for the $i$-th node is defined as
$\text{RWSE}_{i}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d}]_{i,i},$ (6)
Laplacian Positional Encoding. Laplacian positional encoding (LapPE) (Kreuzer et al., 2021; Dwivedi et al., 2023) method emerges as a significant advancement for enriching node representations with spectral information. LapPE utilizes the eigenvectors of the graph Laplacian matrix to encode the relative positions of nodes within a graph. These eigenvectors reflect a local coordinate system that captures meaningful structural information, while also preserving the global topological properties of the original graph. The Laplacian matrix $\mathbf{L}=\mathbf{D}-\mathbf{A}$ has the full eigendecomposition as $\mathbf{L}=\mathbf{U\Lambda U}^{\top}$. We use the $k$ smallest non-trivial eigenvectors of the Laplacian matrix to create the LapPE. The LapPE with $k$ eigenvectors for the $i$-th node is expressed as:
$\text{LapPE}_{i}=\left[\boldsymbol{u}_{1,i},\boldsymbol{u}_{2,i},\cdots,\boldsymbol{u}_{k,i}\right]\in\mathbb{R}^{k},$ (7)
where $\boldsymbol{u}_{t}$ denotes the $t$-th smallest non-trivial eigenvectors and $k$is the number of eigenvectors used. Laplacian PE is particularly useful in situations where nodes are inherently anonymous and lack unique features. However, the Laplacian encoding faces challenges from the arbitrary sign of normalized eigenvectors, introducing $2^{k}$ possible configurations for $k$ eigenvectors. To manage this complexity during training, eigenvectors are randomly sampled from these possibilities. Alternatively, resolving the sign ambiguity by taking the absolute values of eigenvectors simplifies the model but can significantly reduce the expressiveness of the positional features.
Pre-trained Positional and Structural Encoder. Graph Positional and Structural Encoder (GPSE) (Liu et al., 2023b) is a graph encoder pre-trained on molecule datasets by reconstructing traditional positional encoding, such as LapPE, RWSE, CycleSE, etc. The model takes as input graph adjacency matrix and randomly generates node features by $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to improve the expressiveness. The pre-training architecture is a deep MPNN with 20 layers and residual connection and gating mechanism. Moreover, GPSE utilizes a virtual node technique in each graph to enable global message passing. However, GPSE suffers from poor generalizability across other domains, due to its pretraining setting and randomized node features.
Appendix B Dataset Pre-processing
B.1 Shortest Path Distance Regression
Shortest Path Distance (SPD) Regression is an edge-level reconstruction task. SPD encodes the global proximity and connectivity between nodes, which helps to discern nodes’ positions and relations within the entire graph (Li et al., 2020). For data preprocessing, we pre-calculate the $N\times N$ SPD matrix of a given graph before the pretraining phase and save the matrix with the graph data for fast retrieval. We utilize the Dijkstra algorithm (Dijkstra, 1959) to compute the shortest path distances between node pairs, which serve as SPD labels. During pretraining, we randomly select node pairs and fetch their SPD as labels to perform the edge-level reconstruction task. The loss for shortest path distance regression is computed as:
$\mathcal{L}_{\text{SPD}}=\frac{1}{|E|}\sum_{i,j\in V}\|h_{\text{SPD}}(\mathbf{P}^{L}_{i}\|\mathbf{P}^{L}_{j})-\text{SPD}_{i,j}\|^{2}$
[…]
B.2 Motif Counting
Motif Counting is formulated as a node-level reconstruction task. We explicitly use subgraph isomorphism counting as a form of self-supervision during pretraining, allowing the model to better leverage structural information and identify each node’s role in the surrounding subgraphs (Bouritsas et al., 2022). We follow previous works on subgraph isomorphism to count the number of certain motifs surrounding each node. A key concept in our preprocessing step is ”Automorphism orbits,” as introduced by Pržulj (Pržulj, 2007). This concept helps in identifying unique roles of nodes within the counted motifs. The detailed counting method is as follows:
  1. 1.
    Define the subgraph structures to be counted: Enumerate all graphlets with the number of nodes less than or equal to 4. This limitation is due to the exponential growth in the number of graphlets with increasing nodes, which becomes computationally infeasible.
  2. 2.
    Assign indexed orbits: For each subgraph (or motif), assign each vertex a uniquely indexed orbit to facilitate accurate counting.
  3. 3.
    Count specified subgraphs: Utilize the subgraph_isomorphism function from the Python package graph-tool to count specified subgraphs throughout the entire graph. This count is used to determine orbits for each node in the entire graph, forming an orbit degree vector.
  4. 4.
    Save and prepare node-level labels: Save these vectors with the data as node-level labels, which are then prepared for our node-level reconstruction task.
Figure 4: Illustration of various subgraphs (graphlets) used in the motif counting. Each subgraph is indexed and labeled for reference.
Appendix C Methodologies
C.1 Dataset
Table 6: Dataset Information. class. represents classification task and reg. represents regression task.
Dataset Num. Num. Num. Pred. Pred. Num. Metric
graphs nodes edges level task tasks
MolPCBA 437,929 25.97 28.11 graph class. (binary) 128 AP
MolHIV 41,127 25.51 27.46 graph class. (binary) 1 AUROC
MNIST 70,000 70.57 281.65 graph class. (10-way) 1 ACC
Peptides-func 15,535 150.94 153.65 graph class. (binary) 10 AP
Peptides-struct 15,535 150.94 153.65 graph reg. 11 MAE
ogbn-proteins 1 132,534 39,561,252 node class. (binary) 112 AUROC
Pokec 1 1,632,803 30,622,564 node class. (binary) 1 ACC
ogbn-arxiv 1 169,343 1,166,243 node class. (40-way) 1 ACC
ogbn-products 1 2,449,029 61,859,140 node class. (47-way) 1 ACC
ZINC 249,456 23.2 49.8 graph reg. 1 MAE
PubMed 19,717 88,648 500 node class. (3-way) 1 ACC
CIFAR10 60,000 117.6 941.2 graph class. (10-way) 1 ACC
MolPCBA dataset consists of 437,929 molecular graphs, each representing a compound from the PubChem BioAssay database. This dataset is used for multi-label binary classification tasks across 128 targets, focusing on predicting the bioactivity of compounds against various protein targets. The primary evaluation metric for this dataset is Average Precision (AP).
MolHIV contains 41,127 molecular graphs derived from the MoleculeNet benchmark. Each graph represents a molecule, with nodes as atoms and edges as chemical bonds. The task is a binary classification to predict the ability of molecules to inhibit HIV replication, with AUROC as the evaluation metric.
MNIST dataset includes 70,000 images converted into graphs. Each image represents a handwritten digit, with nodes representing pixels and edges representing pixel adjacency. The task is a 10-way classification to identify the digit in the image, evaluated using accuracy (ACC).
Peptides-func comprises 15,535 graphs, each representing a peptide. Nodes represent amino acids, and edges represent peptide bonds. The task involves binary classification to predict the functional properties of the peptides, with Average Precision (AP) as the evaluation metric.
Peptides-struct also contains 15,535 peptide graphs but focuses on regression tasks to predict structural properties of peptides, such as bond angles and distances. The evaluation metric is Mean Absolute Error (MAE).
ogbn-proteins dataset is a large-scale graph with 132,534 nodes and 39,561,252 edges, representing protein-protein interaction networks. Each node is a protein, and edges represent interactions. The task is binary classification at the node level to predict protein functions, evaluated using AUROC.
Pokec is a social network dataset from the Pokec online social network in Slovakia. It includes one large graph with 1,632,803 nodes (users) and 30,622,564 edges (friendships). The task is binary classification to predict user attributes, such as gender, with accuracy (ACC) as the metric.
ogbn-arxiv dataset consists of a single large graph with 169,343 nodes and 1,166,243 edges, representing the citation network of arXiv papers. Each node is a paper, and edges represent citation links. The task is 40-way classification to predict the primary subject area of each paper, evaluated using accuracy (ACC).
ogbn-products dataset includes a large graph with 2,449,029 nodes and 61,859,140 edges, representing an Amazon product co-purchasing network. Nodes represent products, and edges represent co-purchasing relationships. The task is 47-way classification to predict the product category, with accuracy (ACC) as the metric.
ZINC contains 249,456 molecular graphs, where each graph represents a molecule from the ZINC database. The task is regression to predict molecular properties like solubility, with Mean Absolute Error (MAE) as the evaluation metric.
PubMed dataset includes a citation network of 19,717 scientific publications. Nodes represent papers, and edges represent citations. The task is 3-way classification to predict the subject areas of the papers, evaluated using accuracy (ACC).
CIFAR10 dataset has 60,000 images transformed into graphs, where each image represents a colored object. Nodes represent pixels, and edges represent pixel adjacency. The task is a 10-way classification to identify the object in the image, evaluated using accuracy (ACC).
Triangle poses triangle counting as a 10-way graph-level classification task. Half of the test set are graphs with a similar size to those in the training and validation set (denoted as Triangle-S). The left are graphs with larger sizes (denoted as Triangle-L), which present greater challenges to the model’s expressive power.
Pattern and Cluster are graph datasets generated with the Stochastic Block Model (SBM) (Abbe, 2018) that require the model to discern graph patterns and local clusters, which is a node-level classification task.
Table 7: Dataset statistics of three categories from Amazon e-commerce networks
# nodes # edges avg. degree avg. # tokens
Clothing 469,274 2,578,746 10.99 117.83
Home 453,121 3,732,948 16.48 133.94
Sports 293,712 2,390,076 16.27 125.08
C.2 Pre-training Metric
We use an accuracy metric to measure Community Detection and Graph Contrastive Learning and use mean squared error (MSE) to measure the performance of Shortest Path Distance regression and Motif Counting.
  • For the community detection task, we set $\epsilon$ as $1$. The predicted label between the $i$-th node and the $j$-th node $\hat{Y}_{i,j}^{c}$ is $1$ if $\operatorname{sim}(i,j)\geq 0.5$ and 00 otherwise. Accuracy is calculated by comparing the predicted label $\hat{Y}_{i,j}^{c}$ with the ground truth labels $Y_{i,j}^{c}$ and is defined as the proportion of correctly predicted labels out of all possible node pairs:
[…]
    where $\mathbbm{1}(\cdot)$ is an indicator function and $|V|(|V|-1)/2$ is the total number of unique node pairs in the graph. This metric effectively measures how well the model can identify community structures by correctly classifying node pairs as being in the same community or in different communities.
  • For the graph contrastive learning task, we evaluate pre-training performance using the accuracy metric, which measures the model’s ability to correctly classify graphs as originating from the same or different datasets. The accuracy is computed by:
[…]
    where $\hat{Y}_{G_{i},G_{j}}$ is the predicted label indicating whether graph $G_{i}$ and $G_{j}$ are from the same dataset and $Y_{G_{i},G_{j}}$ is the ground truth label. $\hat{Y}_{G_{i},G_{j}}$ is $1$ if $\operatorname{sim}(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}})\geq 0$ and 00 otherwise.$N$ is the total number of evaluated graph pairs.
  • For shortest path distance regression, the mean squared error (MSE) is used as a metric, which is defined as:
[…]
    The ground truth SPD is normalized by the graph diameter to ensure scale consistency and training stability.
  • For the motif counting task, the mean absolute error is used as a metric, which is defined as:
[…]
    where $Y_{i}^{m}$ is the pre-computed label for the $i$-th node.
Table 8: Test performance of the pre-trained model with different architectures on four pre-training tasks.
Architecture Community Detection Graph Contrastive Learning Shortest Path Distance Motif Counting
Accuracy Accuracy MSE MAE (log)
GatedGCN 0.859 0.831 0.155 2.592
GCN 0.883 0.823 0.141 2.593
GIN 0.892 0.853 0.141 2.611
None+TF 0.828 0.841 0.214 2.603
GCN+TF 0.829 0.841 0.212 2.602
GIN+TF 0.829 0.860 0.193 2.603
None+TF(w/ bias) 0.829 0.863 0.218 2.603
GCN+TF(w/ bias) 0.927 0.857 0.134 2.591
GIN+TF(w/ bias) 0.932 0.862 0.114 2.580
C.3 Biased Attention
The illustration of biased attention is shown in Figure 5
Figure 5: Biased Attention based on random walk matrix.
C.4 Uncertainty-based Loss Weighing
The scale of the loss of different tasks can be different, causing the overall loss to be dominated by a certain task, and ultimately the loss of the other tasks cannot affect the learning process of the network-sharing layers. We use the uncertainty-based loss-weighing method (Kendall et al., 2018) to automatically balance the four pre-training tasks and unify the different scales. Moreover, the uncertainty value reflects the contribution of each task towards the overall pre-training process. A higher uncertainty value indicates a lower contribution (Kendall et al., 2018). Let $\sigma_{\tau}$ and $\mathcal{L}_{\tau}$ represent the task-specific uncertainty value for the task $\tau$. The overall pre-training loss is computed by:
$\mathcal{L}=$ $\frac{1}{\sigma_{\text{SPD}}^{2}}\mathcal{L}_{\text{SPD}}+\frac{1}{\sigma_{\text{MC}}^{2}}\mathcal{L}_{\text{MC}}+\frac{1}{\sigma_{\text{CD}}^{2}}\mathcal{L}_{\text{CD}}+\frac{1}{\sigma_{\text{GCL}}^{2}}\mathcal{L}_{\text{GCL}}$ (12)
$+\log\sigma_{\text{SPD}}+\log\sigma_{\text{MC}}+\log\sigma_{\text{CD}}+\log\sigma_{\text{GCL}}.$
C.5 Pre-training Setting
[…]
Appendix D Expressiveness
D.1 Theoretical Proof
For an input graph $G(V,E)$ with node set $V$ and edge set $E$, let $f_{P}:V\rightarrow\mathcal{X}$ and $f_{R}:V\times V\rightarrow\mathcal{X}$ indicate the node-level and edge-level structural encoding, respectively. SEG-WL updates the node labels at the $t$-th iteration by $g_{t}(v)=\text{hash}\left(\left\{\left\{\left(g_{t-1}(u),f_{R}(v,u)\right):u\in V\right\}\right\}\right)$ and $g_{0}(v)=\text{hash}(f_{P}(v))$.
Proposition D.1.
RW($d$)-SEG-WL ($d\geq 3$) is strictly more expressive than WL in testing non-isomorphic graphs.
Proposition D.2.
There exist pairs of graphs that RW($d$)-SEG-WL can distinguish, but 3-WL can not.
Proof.
We first introduce Neighbor-SEG-WL, which is the SEG-WL test when $f_{P}$ is an identity encoding and $f_{R}(u,v)$ equals 1 if $(u,v)\in E$ and $2$ otherwise. Previous works have proved the following Proposition (Zhu et al., 2023).
Proposition D.3.
Two non-isomorphic graphs can be distinguished by WL if and only if they are distinguishable by Neighbor-SEG-WL.
Therefore, Neighbor-SEG-WL is a specific example of SEG-WL test that has equivalent expressiveness to the 1-WL test. We then prove that RW-SEG-WL is strictly more expressive than Neighbor-SEG-WL. Let $d_{\text{neg}}(u,v)$ indicate the edge-level encoding $f_{R}$ in Neighbor-SEG-WL. Note $d_{\text{neg}}(v_{i},v_{j})=2$ if and only if $A_{ij}=0$. Recall that $f_{R}(\cdot,\cdot)$ in RW-SEG-WL satisfies $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ with $\mathbf{R}=[\mathbf{I},\mathbf{M},\cdots,\mathbf{M}^{d-1}]$ where $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$. Therefore, $f_{R}$ in RW-SEG-WL strictly contains the information of $d_{\text{neg}}$. Therefore, if two non-isomorphic graphs can be distinguished by WL, they can be distinguished by RW-SEG-WL. Proposition D.1 is proved.
To prove Proposition D.2, we provide an example in Figure 6 which shows the Shrikhande graph and the Rook’s $4\times 4$ graph, a pair of strongly regular graphs SRG(16,6,2,2). It is proved that they cannot be distinguished by 3-WL (Arvind et al., 2020). We empirically verified that RW($d$)-SEG-WL with $d>4$ can distinguish these two graphs. ∎
Figure 6: RW($d$)-SEG-WL can distinguish the Shrikhande graph and the Rook’s $4\times 4$ graph when $d>4$ while 3-WL fails
D.2 Synthetic Graph Isomorphism Tests
[…]
Table 9: Results of synthetic graph isomorphism tests
Low-Order Graphs (Parameter:$n$) Strongly Regular Graphs (Parameter:($n,k,\lambda,\mu$)
Parameter 5 6 7 8 (25,12,5,6) (26,10,3,4) (29,14,6,7) (36,14,4,6) (40,12,2,4) 45,12,3,3)
# graphs 21 112 853 11117 15 10 41 180 28 78
# graph pairs 210 6216 363378 61788286 105 45 820 16110 378 3003
number of undistinguishable graph pairs
WL 0 3 17 312 105 45 820 16110 378 3003
SPD-SEG-WL 0 2 12 186 105 45 820 16110 378 3003
RW-SEG-WL($d=8$) 0 0 0 0 0 0 0 0 0 0
D.3 Reconstruction of other PSE types
Table 10: Performance of other PSE reconstruction on $5\%$ MolPCBA dataset. The coefficient of determination $R^{2}$ scores are reported as the metric.
PSE type ElstaticPE LapPE RWSE HKdiagSE CycleSE
GPSE 0.964 0.973 0.984 0.981 0.977
Ours 0.947 0.970 0.987 0.984 0.992
Table 10 demonstrates that the PSE generated by our pre-trained GFSE, followed by a trainable lightweight MLP, is capable of reconstructing various pre-defined PSE types on $5\%$ MolPCBA dataset. We evaluate this using the coefficient of determination $R^{2}$ scores as a metric. Notably, our model performs competitively compared with GPSE, achieving an $R^{2}$ of 0.987 for RWSE and 0.992 for CycleSE, given the fact that GPSE directly adopts PSE reconstruction as its training objective. Instead, our method generalizes well across different PSEs without being directly trained for reconstruction. This suggests that our structural self-supervision tasks are effective and sufficient in capturing important structural information.
Appendix E Evaluation Details
E.1 Integration with Pre-trained Models on Molecules
All models are fine-tuned, trained or tested using five different seeds from 42 to 46, with the results averaged. Additionally, for the results of our generated PSE, we select three different seeds to obtain three GFSE checkpoints. Each GFSE is used to run the downstream task five times with the aforementioned seeds ($42-46$), and all results are averaged.
For the training from scratch setting, we adopt and modify the code base from GPS (Rampášek et al., 2022)  [^1] https://github.com/rampasek/GraphGPS. RWSE and LapPE are of dimension $32$ for molecule benchmark in Table 3 across all the datasets. Given a graph $G(V,E)$, we directly concatenate the PSE and the raw node feature as the new input feature, then send them into the very beginning of a model with randomly initialized parameters, which is as follows:
$X^{\prime}=\operatorname{concat}(X,\text{PSE})$ (13)
$\hat{y}=\operatorname{MLP}[\operatorname{pooling}[\operatorname{GraphModel}(X^{\prime})]],$ (14)
where $X$ denotes the raw node feature of the input graph, $X^{\prime}$ is the input feature augmented with structural information, $\operatorname{GraphModel}$ denotes our backbones GNN or GPS, and the read-out layer consists of pooling and MLPs to obtain the final prediction.
For the fine-tuning setting in the molecule benchmark, we concatenate the node encodings obtained from a pre-trained model with the extra structural features. Then we send the concatenated features into the final read-out layers for the final prediction. Note that during the fine-tuning process, the parameters of the entire model (both the pre-trained model and the read-out layer) are continuously updated.
$X^{\prime}=\operatorname{GNN}(X)$ (15)
$\hat{y}=\operatorname{MLP}[\operatorname{pooling}[\operatorname{concact}(X^{\prime},\text{PSE})]],$ (16)
where $X^{\prime}$ denotes the latent node features output from the pre-trained models, PSE denotes the extra structural feature generated by GFSE, the read-out layer consists of pooling and MLPs, where the hyper-parameters follow exactly as  (Xia et al., 2022). We report the performance of the model which achieves the best validation performance during training.
E.2 Integration with Large Language Models
To incorporate the graph structural information into the language model, we use a lightweight MLP to project the $32$-dimensional PSE generated by GFSE into the $4096$-dimensional embedding space of LLaMA2-7B (Touvron et al., 2023). The MLP ensures alignment between the PSE and the language model’s embeddings, facilitating smooth integration of graph information. We then concatenate the textual description of the central node with those of its one-hop neighbors, prepend the projected PSE as a soft token, and append a special graph token at the end of the sequence. This tokenized sequence is fed into the language model, and the hidden embedding corresponding to the special graph token is extracted to represent the central node. We use a contrastive loss between positive node pair $(i,j)\in E$ and negative node pair $(i,j^{\prime})\notin E$ to train the MLP and finetune the language model
$\mathcal{L}=\sum_{(i,j)\in E}\left(d_{ij}^{2}+\max\left(\tau-d_{ij^{\prime}},0\right)^{2}\right),$ (17)
where $d_{ij}=1-\cos(v_{i},v_{j})$ and $v_{i}$ indicates the representation of the node $i$, and $\tau$ is the similarity margin (set as $0.5$ in our experiments).
Appendix F More Experimental Results
F.1 Downstream Evaluation Performance
[…]
ogbn-arxiv PubMed MNIST CIFAR10
GateGCN $71.69_{\pm 0.21}$ $76.86_{\pm 0.41}$ $97.34_{\pm 0.14}$ $67.31_{\pm 0.31}$
GateGCN+LapPE $71.95_{\pm 0.37}$ $74.83_{\pm 0.24}$ $97.10_{\pm 0.28}$ $65.08_{\pm 0.26}$
GateGCN+RWSE $71.83_{\pm 0.65}$ $76.11_{\pm 0.39}$ $96.84_{\pm 0.27}$ $65.26_{\pm 0.68}$
GateGCN+GPSE $72.17_{\pm 0.42}$ $71.97_{\pm 0.36}$ $96.94_{\pm 0.17}$ $65.63_{\pm 0.27}$
GateGCN+GFSE $\mathbf{72.61}_{\pm 0.53}$ $\mathbf{78.39}_{\pm 0.84}$ $\mathbf{97.44}_{\pm 0.31}$ $\mathbf{68.39}_{\pm 0.47}$
Transformer (TF) $5.86_{\pm 0.00}$ $\mathbf{66.63}_{\pm 0.73}$ $97.29_{\pm 0.11}$ $69.04_{\pm 0.28}$
TF+LapPE $5.86_{\pm 0.00}$ $66.27_{\pm 0.46}$ $96.95_{\pm 0.38}$ $69.01_{\pm 0.61}$
TF+RWSE $5.86_{\pm 0.00}$ $64.43_{\pm 0.37}$ $97.81_{\pm 0.58}$ $70.70_{\pm 0.45}$
TF+GPSE $21.56_{\pm 2.74}$ $65.89_{\pm 0.14}$ $97.78_{\pm 0.32}$ $69.57_{\pm 0.16}$
TF+GFSE $\mathbf{23.84}_{\pm 3.15}$ $66.30_{\pm 0.68}$ $\mathbf{98.03}_{\pm 0.84}$ $\mathbf{71.33}_{\pm 0.23}$
GPS $70.68_{\pm 0.71}$ $\mathbf{74.26}_{\pm 0.60}$ $98.05_{\pm 0.12}$ $71.49_{\pm 0.35}$
GPS+LapPE $69.51_{\pm 0.38}$ $73.68_{\pm 0.37}$ $98.16_{\pm 0.28}$ $71.87_{\pm 0.21}$
GPS+RWSE $72.14_{\pm 0.84}$ $72.87_{\pm 0.44}$ $\mathbf{98.19}_{\pm 0.30}$ $71.30_{\pm 0.33}$
GPS+GPSE $71.21_{\pm 0.34}$ $73.71_{\pm 0.70}$ $98.08_{\pm 0.13}$ $72.31_{\pm 0.25}$
GPS+GFSE $\mathbf{72.30}_{\pm 0.13}$ $74.20_{\pm 0.35}$ $98.15_{\pm 0.46}$ $\mathbf{74.11}_{\pm 0.93}$
[…]
Additional evaluation results on ogbn-Arxiv, PubMed, MNIST and CIFAR10 are given in Table 11, where PubMed and CIFAR10 are out of the pre-training dataset. We observe that most structural encodings surprisingly degrade the performance on the PubMed dataset. The proposed GFSE surpasses other PSE on CIFAR10, demonstrating its generalizability in downstream evaluation.
Table 12: Comparison with GraphPEFT
Cloth Home Sport
Hit@1 MRR Hit@1 MRR Hit@1 MRR
GraphPEFT 76.95 84.71 79.87 86.76 64.61 77.34
   w.o. pre-training 76.74 84.57 79.68 86.63 64.44 77.21
LLaMA + GFSE 76.84 84.68 79.85 86.77 64.79 76.24
Table 12 compares our approach with GraphPEFT. GraphPEFT involves two steps: pre-training a graph encoder on domain-specific graph data and fine-tuning it on evaluation datasets. This process requires a domain-specific encoder for each dataset, increasing the adaptation cost when moving across different domains. In contrast, our model is designed for general-domain usage, offering a more flexible and cost-effective adaptation without requiring domain-specific pre-training. As shown in the table, our model consistently outperforms the version of GraphPEFT that skips the domain-specific pre-training step (w.o. pre-training). This further highlights the robustness and generalizability of our approach, as it avoids the need for costly pre-training on specific domains while still achieving competitive or superior results. Specifically, for the “Sport” dataset, our model demonstrates comparable performance to GraphPEFT with pre-training, further underscoring the adaptability of GFSE in varied contexts.
F.2 Efficiency Evaluation
During the pre-training stage on eight datasets, the average time is around 30 to 40 minutes for each epoch with a single NVIDIA A40 48GB GPU. Total training time is less than two days, which is relatively efficient for a comprehensive multi-dataset pre-training process.
Table 13: Runntimes ($s$) of PSE computation on random synthetic graph with increasing numbers of nodes
PSE / Graph size 100 300 500 1000
LapPE 2 9.25 34 155
RWSE 2 9.76 31.48 207
Pre-computation 0.0007 0.001 0.003 0.006
GFSE Inference 0.908 3.958 10.770 48.106
We compare the inference efficiency of GFSE with handcrafted positional encodings, such as LapPE and RWSE in Table 13 and Table 14. Specifically, we generate 1,000 synthetic Erdos-Rényi graphs for various graph sizes (100, 300, 500, and 1,000 nodes) and evaluate the time required for pre-computation and inference in GFSE.
Table 14: Runntimes ($s$) on real-world graph dataset
Dataset ZINC MolHIV MolPCBA Peptides MNIST CIFAR10
LapPE 25 sec 37 sec 6.13 min 73 sec 96 sec 2.55 min
RWSE 11 sec 58 sec 8.33 min - - -
GFSE 4.17 sec 17.23 sec 2.97 min 15.21 sec 49.38 sec 1.27 min
As shown in the table, both LapPE and RWSE exhibit significant increases in computation time as the graph size grows. Pre-computation times required by GFSE inference remain minimal for all graph sizes, underlining the model’s efficiency in this phase. In Table 14, we observe that GFSE demonstrates superior scalability in inference, making it a more efficient option for large-scale graph processing.
Appendix G Discussion
Limitation and Social Impact. While GFSE represents a significant step forward in developing general and expressive foundation models for graph-structured data, there are certain limitations to consider. For example, the effectiveness of GFSE may be influenced by the quality of the pre-training datasets, as biases or under-representation in the data could propagate to the learned representations. Future work could focus on expanding the diversity of pre-training datasets to mitigate such biases and improve the robustness of GFSE across more diverse domains
From a social impact perspective, the development of structural graph encoder like GFSE holds significant promise in advancing various application domains that rely on graph analytics. In fields such as computational biology, social network analysis, and recommendation systems, GFSE could enable more accurate and efficient modeling of complex structured data, leading to improved understanding and decision-making. Furthermore, by reducing the need for extensive task-specific fine-tuning, GFSE could democratize the use of powerful graph learning techniques, making them more accessible to researchers and practitioners with limited computational resources.